Peter Parker is evaluating two different methods for manufacturing spiderwebs.

The first option is manufacturing them in his home lab. This will entail purchasing an industrial-grade chemical reactor that costs $450,000. He will have to hire a lab assistant who will initially receive a salary of $60,000 a year, but that salary must be adjusted for inflation every year at an inflation rate of 2.5%. After ten years, the chemical reactor can be sold for $30,000. This lab setup can produce 10,000 cubic meters of spiderwebs per year.

The second option is outsourcing all the manufacturing process to a biotech company recently set up by Bruce Banner. In this case, the price of the spiderwebs will be $14 per cubic meter over the next ten years.

Since Peter Parker is a broke student, all of this project is being financed by a superhero fund set up by his late friend and mentor Tony Stark. This fund returns 11% annually. Since this is not a commercial endeavor, there are no taxes nor depreciation to consider. Given these data, calculate the unit price of manufacturing spiderwebs in the lab and decide which of the two options makes the most economic sense. Justify your answers with calculations.
Answer #1

Present worth of geometric series = A*[1 - (1+g)^n / (1+i)^n]/(i-g)

Present cost of salary = 60000*[1 - (1+0.025)^10 / (1+0.11)^10]/(0.11-0.025)

= 60000*[1 - (1.025)^10 / (1.11)^10]/(0.085)

= 60000* 6.46087167

= 387652.30

EUAC of manufacturing 10000 webs = 450000*(A/P,11%,10) + 387652.30 *(A/P,11%,10) - 30000*(A/F,11%,10)

= 450000*0.169801 + 387652.30 *0.169801 - 30000*0.059801

= 140440.17

Manufacturing Cost per unit of web = 140440.17 / 10000 = 14.04

Outsourcing Cost per unit = 14

As Outsourcing Cost per unit is less, it should be selected